# The free energy of Dirac’s vacuum in purely magnetic fields

### v1

@@ 2. Derivation of the model and main results
The function $f_{\mathrm{PV}}^{0}$ is the well known Euler-Heisenberg energy function at null temperature, which has been introduced for the first time in [37]. Actually, the authors here derive a concave-decreasing and negative function given by (2) and diverging faster than $\left\lvert B\right\rvert^{2}$ at infinity. For this reason, the total energy of the system, given in the limit by an effective local Lagrangian, is unbounded from below and the model is unstable. Then, the same function has been mathematically studied in the same regime of slowly varying purely magnetic fields in [31], where the authors derive a non-negative Euler-Heisenberg energy function which guarantees the corresponding model to be stable. Moreover, they find out that the instability of the previous Lagrangian was due to the charge renormalisation [31, Section 2.3]. On the other hand, the function $f_{\mathrm{PV}}^{T}$ is the non-zero temperature contribution to the total Euler-Heisenberg energy function. To our knowledge, this is the first rigorous derivation of this term, which can be found in physics literature in [16] and [17, Equation (3.162)]. Notice that actually Eq. 19 is slightly different from Eq. 5 appearing in [16]. Indeed, the thermal Euler-Heisenberg formula found in the physics literature is not a well defined object because of the well known ultraviolet divergence in QED. As previously explained, here we deal with this issue by using the Pauli-Villars regularisation technique. In our framework, it correponds to the appearance of the sum in $j$ and consequently of the masses $m_{j}$’s. Moreover, the cotangent in (5) makes the integral obviously ill-defined due to its infinitely many poles. The concurrent presence of the imaginary exponential suggests to rotate the integration path in order to replace the cotangent by the hyperbolic cotangent by getting a formula closer to ours. This procedure is standard in the physics literature, but of course should be carefully justified from a mathematical point of view. The proof of 2.3 can be found in Section 4.
@@ 3. Rigorous definition of the PV-regularised free energy functional
When $S$ is a self-adjoint operator on $L^{2}\left(\mathbb{R}^{3},\mathbb{R}^{4}\right)$, with domain $D\left(S\right)$, it follows from (21) using standard functional calculus [46] that the hyberbolic tangent $\tanh{S}$ of $S$ is given by
@@
$\left\lVert\frac{4S^{2}}{4S^{2}+\left(2k-1\right)^{2}\pi^{2}}\right\rVert_{D\left(S^{2}\right)\xrightarrow{}L^{2}\left(\mathbb{R}^{3},\mathbb{R}^{4}\right)}\leq\min\{1,\left(\left(2k-1\right)\pi\right)^{-2}\left\lVert 4S^{2}\right\rVert_{D\left(S^{2}\right)\xrightarrow{}L^{2}\left(\mathbb{R}^{3},\mathbb{R}^{4}\right)}\}.$
@@
In order to prove the theorem, we want to prove that the $\mathbb{C}^{4}-$trace of the series in the right-hand side of (22) defines a trace class operator whose Schatten norm can be estimated according to the following equality:
$\sum_{k\in\mathbb{Z}}\left\lVert\sum_{j=0}^{2}c_{j}\operatorname{tr}_{\mathbb{C}^{4}}\left(\frac{i\omega\left(k,\beta\right)}{D^{A}_{m_{j}}+i\omega\left(k,\beta\right)}-\frac{i\omega\left(k,\beta\right)}{D^{A}_{m_{j}}-i\omega\left(k,\beta\right)}\right.\right.\\ \left.\left.-\frac{i\omega\left(k,\beta\right)}{D^{0}_{m_{j}}+i\omega\left(k,\beta\right)}+\frac{i\omega\left(k,\beta\right)}{D^{0}_{m_{j}}-i\omega\left(k,\beta\right)}\right)\right\rVert_{\mathfrak{S}_{1}}=O\mathopen{}\left(\beta\right),$
as $\beta\xrightarrow{}0$, which can be proved when $A\in L^{1}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)\cap H^{1}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)$.
We just need to derive estimates in Schatten spaces of the operator
$\mathcal{R}\left(\omega\left(k,\beta\right),A\right)=\sum_{j=0}^{2}c_{j}\operatorname{tr}_{\mathbb{C}^{4}}\left(\frac{i\omega\left(k,\beta\right)}{D^{A}_{m_{j}}+i\omega\left(k,\beta\right)}-\frac{i\omega\left(k,\beta\right)}{D^{A}_{m_{j}}-i\omega\left(k,\beta\right)}\right.\\ \left.-\frac{i\omega\left(k,\beta\right)}{D^{0}_{m_{j}}+i\omega\left(k,\beta\right)}+\frac{i\omega\left(k,\beta\right)}{D^{0}_{m_{j}}-i\omega\left(k,\beta\right)}\right),$
@@
$\mathcal{R}\left(\omega\left(k,\beta\right),A\right)=\sum_{n=1}^{5}\operatorname{tr}_{\mathbb{C}^{4}}\left(R_{n}\left(\omega\left(k,\beta\right),A\right)+R_{n}\left(-\omega\left(k,\beta\right),A\right)\right)\\ +\operatorname{tr}_{\mathbb{C}^{4}}\left(R^{\prime}_{6}\left(\omega\left(k,\beta\right),A\right)+R^{\prime}_{6}\left(-\omega\left(k,\beta\right),A\right)\right),$
@@
Basically, by replacing the integral with respect to $\omega$ with the series with respect to $k$, the same proof as in [28, Theorem 2.1] holds giving that
$\sum_{k\in\mathbb{Z}}\left(\sum_{j=1}^{5}\left\lVert\operatorname{tr}_{\mathbb{C}^{4}}\left(R_{n}\left(\omega\left(k,\beta\right),A\right)+R_{n}\left(-\omega\left(k,\beta\right),A\right)\right)\right\rVert_{\mathfrak{S}_{1}}\right.\\ \left.+\left\lVert\operatorname{tr}_{\mathbb{C}^{4}}\left(R^{\prime}_{6}\left(\omega\left(k,\beta\right),A\right)+R^{\prime}_{6}\left(-\omega\left(k,\beta\right),A\right)\right)\right\rVert_{\mathfrak{S}_{1}}\right)=O\mathopen{}\left(\beta\right),$
as $\beta\xrightarrow{}0$.∎
@@ 3.2. Gauge-invariant estimates
$\operatorname{tr}_{\mathbb{C}^{4}}T_{A}\left(\beta\right)=\frac{1}{2}\sum_{j=0}^{2}c_{j}\operatorname{tr}_{\mathbb{C}^{4}}\left(D^{0}_{m_{j}}\tanh{\frac{\beta D^{0}_{m_{j}}}{2}}-D^{A}_{m_{j}}\tanh{\frac{\beta D^{A}_{m_{j}}}{2}}\right)$
is trace class when the magnetic potential $A$ decays fast enough, namely when $A\in L^{1}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)\cap\dot{H}^{1}_{\mathrm{div}}\left(\mathbb{R}^{3}\right)$. More precisely, in the proof of the first part of 2.2, we write
@@
with $R_{n}$, $R^{\prime}_{6}$ given by (23) and (24), and we prove that the operators $\operatorname{tr}_{\mathbb{C}^{4}}T_{n}\left(A,\beta\right)$ and $\operatorname{tr}_{\mathbb{C}^{4}}T^{\prime}_{6}\left(A,\beta\right)$ are trace class. Nevertheless, our estimates involve non gauge-invariant quantities (some $L^{p}$ norms of $A$) and they require that $A$ decays fast enough at infinity. We are interested in estimates which involve only the norm $\left\lVert\mathrm{curl}A\right\rVert_{L^{2}}$.
First of all, let us state the following lemma which says that the trace of the odd order operators $\operatorname{tr}_{\mathbb{C}^{4}}T_{1}\left(A,\beta\right)$, $\operatorname{tr}_{\mathbb{C}^{4}}T_{3}\left(A,\beta\right)$ and $\operatorname{tr}_{\mathbb{C}^{4}}T_{5}\left(A,\beta\right)$ vanishes. This is a consequence of the charge-conjugation invariance which is usually known as Furry’s theorem in the physics literature (see [25] and [32, Sec. 4.1]).
@@
For $A\in L^{1}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)\cap\dot{H}^{1}_{\mathrm{div}}\left(\mathbb{R}^{3}\right)$ and $n=1,3,5$, we have
$\frac{1}{\beta}\int_{0}^{\beta}\operatorname{tr}\left(\operatorname{tr}_{\mathbb{C}^{4}}T_{n}\left(A,b\right)\right)\mathop{}\mathopen{}\mathrm{d}b\\ =\frac{1}{2\beta}\int_{0}^{\beta}\sum_{k\in\mathbb{Z}}\operatorname{tr}\left(\operatorname{tr}_{\mathbb{C}^{4}}\left(R_{n}\left(\omega\left(k,b\right),A\right)+R_{n}\left(-\omega\left(k,b\right),A\right)\right)\right)\frac{\mathop{}\mathopen{}\mathrm{d}b}{b}=0.$
@@
The proof is the same as in [28, Lemma 4.1]. ∎
We now compute exactly the second order term $\mathcal{F}_{\mathrm{PV},2}$ associated to the term $T_{2}\left(A,\beta\right)$ appearing in the decomposition of $T_{A}\left(\beta\right)$, assuming that $A\in H^{1}\left(\mathbb{R}^{3}\right)$ and $\mathrm{div}A=0$.
@@
$\frac{1}{\beta}\int_{0}^{\beta}\operatorname{tr}\left(\operatorname{tr}_{\mathbb{C}^{4}}T_{2}\left(A,b\right)\right)\mathop{}\mathopen{}\mathrm{d}b=\frac{1}{8\pi}\int_{\mathbb{R}^{3}}\left(M^{0}\left(q\right)+M^{T}\left(q,\beta\right)\right)\left\lvert\widehat{B}\left(q\right)\right\rvert^{2}\mathop{}\mathopen{}\mathrm{d}q=\mathcal{F}_{\mathrm{PV},2}\left(B,\beta\right),$ (25)
@@
The function $M^{0}$ in formula (25) is the dielectric response of Dirac’s vacuum at zero temperature which has already been rigorously computed in [28, Formula (2.21)] and is well known in physics literature [32, Eq. (5.39)]. On the other hand, $M^{T}$ defined in (27) is the contribution at finite temperature $T$ to the dielectric response of Dirac’s vacuum. Notice that the integral with respect to $t$ in (27) is perfectly defined and converges for $q=0$, namely $M^{T}$ is bounded in a neighborhood of $q=0$. This means that we cannot expect an effect similar to the one shown in [33] for the reduced BDF model, where the authors prove that due to the positive temperature the particles in the polarised Dirac sea rearrange themselves so as to completely screen the external potential. As already mentioned, this phenomenon is known as Debye screening in non-relativistic fermionic plasma physics. After all, this could be explained by the nature of the magnetic field which basically acts on charged particles by confining them in a space region rather than attracting or repelling them.
@@
First of all, in order to include a finite temperature notice that we simply have to make the following formal replacements:
@@
where $\beta=1/T\in\mathopen{(}0\mathclose{}\mathpunct{},\infty\mathclose{)}$, and then to consider the averaged primitive function with respect to $\beta$ of each term in the expansion. By keeping this prescription in mind, the same computations as in the proof of [28, Lemma 4.2] yields
$\operatorname{tr}\left(\operatorname{tr}_{\mathbb{C}^{4}}T_{2}\left(A,\beta\right)\right)=\int_{\mathbb{R}^{3}}G\left(q,\beta\right)\left\lvert\widehat{A}\left(q\right)\right\rvert^{2}\mathop{}\mathopen{}\mathrm{d}q,$
@@
We then use the identity (see [32, Chap. 5])
@@
$G\left(q,\beta\right)=\frac{1}{\beta\pi^{3}}\int_{0}^{1}\int_{0}^{\infty}\sum_{j=0}^{2}c_{j}e^{-s\left(m_{j}^{2}+\left(1-u\right)q^{2}\right)}s^{2}\mathop{}\mathopen{}\mathrm{d}s\,u\mathop{}\mathopen{}\mathrm{d}u\int_{\mathbb{R}^{3}}\left(p\cdot q\right)e^{-s\left(p^{2}-2\left(1-u\right)p\cdot q\right)}\mathop{}\mathopen{}\mathrm{d}p\\ \times\left(\sum_{k\in\mathbb{Z}}e^{-s\omega\left(k,\beta\right)^{2}}\omega\left(k,\beta\right)^{2}\right).$
@@
By integrating by parts, we then have
@@
where in the second equality we first use Poisson summation formula [3, Formula 6.5] and then extract the $n=0$ term. Thus, our result can be expressed so far as
@@
contains the non-zero temperature contribution. Let us now focus on the function $G^{0}$. By integrating by parts and again thanks to conditions (12), we have
@@
We then use the identity (see [27, Formulas (3.324(1)) and (3.471(9))])
@@
Now, by using the Sommerfeld integral representation [44, pp. 328] for Bessel functions of second kind,
@@
Notice that by (30) it is easy to see that $G^{T}$ goes to zero as $T$ goes to zero, or equivalently $\beta$ goes to infinity, so as to recover the well known formula for the dielectric response of Dirac’s vacuum at null temperature [28, Lemma 4.2]. ∎
@@
In our framework, by doing the same calculations with an electric potential $V$, we get a different function $G$ since the beginning. By looking at the limit of this function as $\beta$ goes to infinity and replacing the Riemann sum by an integral, we recover the same result found in [28, Formula (4.16)] in the null temperature case. ∎
@@
The function $M^{T}$ given by (27) is well defined, bounded and such that
@@
$\frac{e^{-X_{j}\left(b,u,q\right)\cosh t}\left(1+X_{j}\left(b,u,q\right)\cosh t+e^{-X_{j}\left(b,u,q\right)\cosh t}\right)}{\left(1+e^{-X_{j}\left(b,u,q\right)\cosh t}\right)^{2}}\\ \leq e^{-X_{j}\left(b,u,q\right)\cosh t}\left(2+X_{j}\left(b,u,q\right)\cosh t\right)\\ \leq 2\left(e^{-X_{j}\left(b,u,q\right)\cosh t}+e^{-\frac{1}{2}X_{j}\left(b,u,q\right)\cosh t}\right)\\ \leq 4e^{-\frac{1}{2}X_{j}\left(b,u,q\right)\cosh t}\\ \leq 4e^{-\frac{1}{2}X_{j}\left(b,u,q\right)\left(1+\frac{t^{2}}{2}\right)},$
@@
$\frac{4}{\beta}\int_{0}^{1}u\left(1-u\right)\int_{0}^{\beta}e^{-\frac{1}{2}X_{j}\left(b,u,q\right)}\int_{0}^{\infty}e^{-\frac{t^{2}}{4}X_{j}\left(b,u,q\right)}\mathop{}\mathopen{}\mathrm{d}t\mathop{}\mathopen{}\mathrm{d}b\mathop{}\mathopen{}\mathrm{d}u\\ =\frac{2\sqrt{\pi}}{3\sqrt{m_{j}}\beta}\int_{0}^{\beta}\frac{1}{\sqrt{X_{j}\left(b,u,q\right)}}e^{-\frac{1}{2}X_{j}\left(b,u,q\right)}\mathop{}\mathopen{}\mathrm{d}b\\ \leq\frac{4\sqrt{\pi}}{3\sqrt{m_{j}}}\frac{1}{\sqrt{\beta}}.$
In particular, the previous calculations show that the function $M^{T}$ is well defined and bounded. Now, in order to conclude the proof, let us recall that
@@
and $G\left(q,\beta\right)\geq 0$ by [31, Lemma 11]. Indeed,

### v2

@@ 2. Derivation of the model and main results
$f^{T}_{\mathrm{PV}}\left(a,\beta\right)=\frac{1}{4\pi^{2}}\int_{0}^{\infty}\sum_{j=0}^{2}c_{j}e^{-sm_{j}^{2}}\left(sa\coth\left(sa\right)-1\right)\sum_{n=1}^{\infty}\left(-1\right)^{n}e^{-\frac{\beta^{2}n^{2}}{4s}}\frac{\mathop{}\mathopen{}\mathrm{d}s}{s^{3}}.$ (19)
The additional regularity needed for this result is expected to be completely technical, as explained in [34, Remark 3]. The function $f_{\mathrm{PV}}^{0}$ is the well known Euler-Heisenberg energy function at null temperature, which has been introduced for the first time in [40]. Actually, the authors there derive a concave-decreasing and negative function given by (2) and diverging faster than $\left\lvert B\right\rvert^{2}$ at infinity. For this reason, the total energy of the system, given in the limit by an effective local Lagrangian, is unbounded from below and the model is unstable. Then the same function has been mathematically studied in the same regime of slowly varying purely magnetic fields in [34], where the authors derive a non-negative Euler-Heisenberg energy function which guarantees the corresponding model to be stable. Moreover, they find out that the instability of the previous Lagrangian was due to the charge renormalisation [34, Section 2.3]. On the other hand, the function $f_{\mathrm{PV}}^{T}$ is the non-zero temperature contribution to the total Euler-Heisenberg energy function. To our knowledge, this is the first rigorous derivation of this term, which can be found in the physics literature in [17] and [18, Equation (3.162)].
Remark 2.4.
Notice that the thermal contributions in 2.2 and 2.3 (respectively, $M^{T}$ and $f^{T}_{\mathrm{PV}}$) are convergent in the ultraviolet limit (that is, when $m_{1},m_{2}$ go to +$\infty$+ ∞), meaning that none of these terms is divergent in the high-frequencies domain. This suggests that the well known UV divergences are completely independent on the temperature and arise from the zero temperature theory. In particular, the final results of this paper could be made independent on the regularisation parameters by means of the same renormalisation procedure studied at null temperature [34, Section 2.3]. ∎
Notice also that Eq. 19 is slightly different from Eq. 5 appearing in [17]. Indeed, the thermal Euler-Heisenberg formula found in the physics literature is not a well defined object because of the well known ultraviolet divergences in QED. As previously explained, here we deal with this issue by using the Pauli-Villars regularisation technique. In our framework, it correponds to the appearance of the sum in $j$ and consequently of the masses $m_{j}$’s. Moreover, the cotangent in (5) makes the integral obviously ill-defined due to its infinitely many poles. Thus, we reckon that Dittrich’s result should be interpreted as follows: after taking an integration path which contours the poles in order to avoid them, the presence of the imaginary exponential in (5) suggests to rotate this integration domain, leading to the replacement of the cotangent by the hyperbolic cotangent and getting a formula closer to ours in the UV limit. This procedure is standard in the physics literature [51, 20], but of course should be carefully justified from a mathematical point of view. The proof of 2.3 can be found in Section 4.
@@ 3. Rigorous definition of the PV-regularised free energy functional
When $S$ is a self-adjoint operator on $L^{2}\left(\mathbb{R}^{3},\mathbb{R}^{4}\right)$, with domain $D\left(S\right)$, it follows from (21) using standard functional calculus [49] that the hyberbolic tangent $\tanh{S}$ of $S$ satisfies
@@
$\left\lVert\frac{4S^{2}}{4S^{2}+\left(2k-1\right)^{2}\pi^{2}}\right\rVert_{D\left(S^{2}\right)\xrightarrow{}L^{2}\left(\mathbb{R}^{3},\mathbb{R}^{4}\right)}\leq\min\left\{1,\left(\left(2k-1\right)\pi\right)^{-2}\left\lVert 4S^{2}\right\rVert_{D\left(S^{2}\right)\xrightarrow{}L^{2}\left(\mathbb{R}^{3},\mathbb{R}^{4}\right)}\right\}.$
@@
In order to prove the theorem, we want to prove that the the series in the right-hand side of (22) defines a trace class operator whose Schatten norm can be estimated as follows:
$\sum_{k\in\mathbb{Z}}\left\lVert\sum_{j=0}^{2}c_{j}\left(\frac{i\omega\left(k,\beta\right)}{D^{A}_{m_{j}}+i\omega\left(k,\beta\right)}-\frac{i\omega\left(k,\beta\right)}{D^{A}_{m_{j}}-i\omega\left(k,\beta\right)}\right.\right.\\ \left.\left.-\frac{i\omega\left(k,\beta\right)}{D^{0}_{m_{j}}+i\omega\left(k,\beta\right)}+\frac{i\omega\left(k,\beta\right)}{D^{0}_{m_{j}}-i\omega\left(k,\beta\right)}\right)\right\rVert_{\mathfrak{S}_{1}}=O\mathopen{}\left(\beta\right),$
as $\beta\xrightarrow{}0$, which can be proved when $A\in L^{1}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)\cap H^{1}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)$. Indeed, it is not important that $\mathrm{div}\,A=0$ for this result and consequently we do not require $A$ to be in $\dot{H}^{1}_{\mathrm{div}}\left(\mathbb{R}^{3}\right)$.
In order to do so, we just need to derive estimates in Schatten spaces of the operator
$\mathcal{R}\left(\omega\left(k,\beta\right),A\right)=\sum_{j=0}^{2}c_{j}\left(\frac{i\omega\left(k,\beta\right)}{D^{A}_{m_{j}}+i\omega\left(k,\beta\right)}-\frac{i\omega\left(k,\beta\right)}{D^{A}_{m_{j}}-i\omega\left(k,\beta\right)}\right.\\ \left.-\frac{i\omega\left(k,\beta\right)}{D^{0}_{m_{j}}+i\omega\left(k,\beta\right)}+\frac{i\omega\left(k,\beta\right)}{D^{0}_{m_{j}}-i\omega\left(k,\beta\right)}\right),$
@@
$\mathcal{R}\left(\omega\left(k,\beta\right),A\right)=\sum_{n=1}^{5}\left(R_{n}\left(\omega\left(k,\beta\right),A\right)+R_{n}\left(-\omega\left(k,\beta\right),A\right)\right)\\ +\left(R^{\prime}_{6}\left(\omega\left(k,\beta\right),A\right)+R^{\prime}_{6}\left(-\omega\left(k,\beta\right),A\right)\right),$
@@
Basically, by replacing the integral with respect to $\omega$ by the series with respect to $k$, the same proof as in [30, Theorem 2.1] holds giving that
$\sum_{k\in\mathbb{Z}}\left(\sum_{j=1}^{5}\left\lVert R_{n}\left(\omega\left(k,\beta\right),A\right)+R_{n}\left(-\omega\left(k,\beta\right),A\right)\right\rVert_{\mathfrak{S}_{1}}\right.\\ \left.+\left\lVert R^{\prime}_{6}\left(\omega\left(k,\beta\right),A\right)+R^{\prime}_{6}\left(-\omega\left(k,\beta\right),A\right)\right\rVert_{\mathfrak{S}_{1}}\right)=O\mathopen{}\left(\beta\right),$
as $\beta\xrightarrow{}0$. In particular, the dependence of $O\mathopen{}\left(\beta\right)$ on the magnetic potential $A$ occurs through the $L^{p}$-norms of it, for $p\in\mathopen{\llbracket}1\mathclose{}\mathpunct{},5\mathclose{\rrbracket}$, as well as the $L^{2}$-norm of its associated magnetic field. Therefore, these estimates are not gauge-invariant and this fact justifies the following subsection.∎
Remark 3.1.
Notice that at this stage it could have been possible to include an external electric potential as well. Indeed, the same estimates as the ones in [30, Proof of Proposition 3.1] hold as long as the $\mathbb{C}^{4}$-trace is applied before taking the trace over $L^{2}\left(\mathbb{R}^{3},\mathbb{C}\right)$ (see the comment just after [30, Theorem 2.1] for further details). Nevertheless, we are not able to prove the results in the following subsection in the complete electromagnetic case, mainly because of the lack of the integration by parts due to the discretisation of $\omega$ (integrals in $d\omega$ are replaced by series over a discrete set of $\omega$ at finite temperature). Moreover, a rigorous derivation of the Euler-Heisenberg formula in presence of an electric field is a really complicated task for the moment and exists not even at null temperature. For all of these reasons, we decided to restrict this paper to the purely magnetic case since the beginning, as mentioned in the introduction. ∎
@@ 3.2. Gauge-invariant estimates
$T_{A}\left(\beta\right)=\frac{1}{2}\sum_{j=0}^{2}c_{j}\operatorname{tr}_{\mathbb{C}^{4}}\left(D^{0}_{m_{j}}\tanh{\frac{\beta D^{0}_{m_{j}}}{2}}-D^{A}_{m_{j}}\tanh{\frac{\beta D^{A}_{m_{j}}}{2}}\right)$
is trace class on $L^{2}\left(\mathbb{R}^{3},\mathbb{C}^{4}\right)$ when the magnetic potential $A$ decays fast enough, namely when $A\in L^{1}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)\cap H^{1}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)$. More precisely, in the proof of the first part of 2.2, we write
@@
with $R_{n}$, $R^{\prime}_{6}$ given by (23) and (24), and we prove that the operators $T_{n}\left(A,\beta\right)$ and $T^{\prime}_{6}\left(A,\beta\right)$ are trace class. Nevertheless, our estimates involve non gauge-invariant quantities (some $L^{p}$ norms of $A$) and they require that $A$ decays fast enough at infinity. We are interested in estimates which involve only the norm $\left\lVert\mathrm{curl}A\right\rVert_{L^{2}}$. Notice that in the previous section it was not essential that $\mathrm{div}A=0$. We will need this hypothesis in the sequel of this work (see 3.3 below).
First of all, let us state the following lemma which says that the trace of the odd order operators $T_{1}\left(A,\beta\right)$, $T_{3}\left(A,\beta\right)$ and $T_{5}\left(A,\beta\right)$ vanishes. This is a consequence of the charge-conjugation invariance which is usually known as Furry’s theorem in the physics literature (see [27] and [35, Sec. 4.1]).
@@
For $A\in L^{1}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)\cap H^{1}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)$ and $n=1,3,5$, we have
$\frac{1}{\beta}\int_{0}^{\beta}\operatorname{tr}\left[T_{n}\left(A,b\right)\right]\mathop{}\mathopen{}\mathrm{d}b\\ =\frac{1}{2\beta}\int_{0}^{\beta}\sum_{k\in\mathbb{Z}}\operatorname{tr}\left[R_{n}\left(\omega\left(k,b\right),A\right)+R_{n}\left(-\omega\left(k,b\right),A\right)\right]\frac{\mathop{}\mathopen{}\mathrm{d}b}{b}=0.$
@@
The proof is the same as in [30, Lemma 4.1]. ∎
We now compute exactly the second order term $\mathcal{F}_{\mathrm{PV},2}$ associated to the term $T_{2}\left(A,\beta\right)$ appearing in the decomposition of $T_{A}\left(\beta\right)$, assuming that $A\in H^{1}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)$ and $\mathrm{div}A=0$.
@@
$\frac{1}{\beta}\int_{0}^{\beta}\operatorname{tr}\left[T_{2}\left(A,b\right)\right]\mathop{}\mathopen{}\mathrm{d}b=\frac{1}{8\pi}\int_{\mathbb{R}^{3}}\left(M^{0}\left(q\right)+M^{T}\left(q,\beta\right)\right)\left\lvert\widehat{B}\left(q\right)\right\rvert^{2}\mathop{}\mathopen{}\mathrm{d}q=\mathcal{F}_{\mathrm{PV},2}\left(B,\beta\right),$ (25)
@@
The function $M^{0}$ in formula (25) is the dielectric response of Dirac’s vacuum at zero temperature which has already been rigorously computed in [30, Formula (2.21)] and is well known in the physics literature [35, Eq. (5.39)]. On the other hand, $M^{T}$ defined in (27) is the contribution at finite temperature $T$ to the dielectric response of Dirac’s vacuum. Notice that the integral with respect to $t$ in (27) is perfectly defined and converges for $q=0$, namely $M^{T}$ is bounded in a neighborhood of $q=0$. This means that we cannot expect an effect similar to the one shown in [36] for the reduced BDF model, where the authors prove that due to the positive temperature the particles in the polarised Dirac sea rearrange themselves so as to completely screen the external potential. As already mentioned, this phenomenon is known as Debye screening in non-relativistic fermionic plasma physics. After all, this could be explained by the nature of the magnetic field which basically acts on charged particles by confining them in a space region rather than attracting or repelling them.
@@
First of all, by comparing the calculations in Section 3.1 to the calculations at zero temperature [31], notice that we simply have to make the following formal replacements in order to include a finite temperature:
@@
where $\beta=1/T\in\mathopen{(}0\mathclose{}\mathpunct{},\infty\mathclose{)}$, and then to consider the averaged primitive function with respect to $\beta$ of each term in the expansion. By keeping this prescription in mind, the same computations as in the proof of [30, Lemma 4.2] yields
$\operatorname{tr}\left[T_{2}\left(A,\beta\right)\right]=\int_{\mathbb{R}^{3}}G\left(q,\beta\right)\left\lvert\widehat{A}\left(q\right)\right\rvert^{2}\mathop{}\mathopen{}\mathrm{d}q,$
@@
We then use the identity (see [35, Chap. 5])
@@
$G\left(q,\beta\right)=\frac{1}{\beta\pi^{3}}\int_{0}^{1}\int_{0}^{\infty}\sum_{j=0}^{2}c_{j}e^{-s\left(m_{j}^{2}+\left(1-u\right)q^{2}\right)}\left(\sum_{k\in\mathbb{Z}}e^{-s\omega\left(k,\beta\right)^{2}}\omega\left(k,\beta\right)^{2}\right)\\ \times\int_{\mathbb{R}^{3}}\left(p\cdot q\right)e^{-s\left(p^{2}-2\left(1-u\right)p\cdot q\right)}\mathop{}\mathopen{}\mathrm{d}p\,s^{2}\mathop{}\mathopen{}\mathrm{d}s\,u\mathop{}\mathopen{}\mathrm{d}u.$
@@
Integrating by parts, we then get
@@
where in the second equality we first use Poisson summation formula [4, Formula 6.5] and then extract the $n=0$ term. Thus, our result can be expressed so far as
@@
contains the non-zero temperature contribution. Let us now focus on the function $G^{0}$. Integrating by parts and again thanks to conditions (12), we have
@@
We then use the identity (see [29, Formulas (3.324(1)) and (3.471(9))])
@@
Now, by exploiting the Sommerfeld integral representation [47, pp. 328] for Bessel functions of second kind,
@@
The formal replacement presented at the beginning of the proof is strictly related to the KMS condition which prescribes the anti-periodicity with respect to an imaginary time for Green’s functions. However, this substitution is derived here by following a completely different argument, which is purely mathematical. More specifically, the replacement of integrals over frequencies by series basically arises from the representation of $x\tanh x$ as a power series. ∎
@@
Notice that by (30) it is easy to see that $G^{T}$ goes to zero as $T$ goes to zero, or equivalently $\beta$ goes to infinity, so as to recover the well known formula for the dielectric response of Dirac’s vacuum at null temperature [30, Lemma 4.2]. ∎
Remark 3.6.
In our framework, by doing the same calculations with an electric potential $V$, we get a different function $G$ since the beginning. By looking at the limit of this function as $\beta$ goes to infinity and replacing the Riemann sum by an integral, we recover the same result found in [30, Formula (4.16)] in the null temperature case. ∎
@@
The function $M^{T}$ given by (27) is well defined, bounded for any fixed $\beta\in\mathopen{(}0\mathclose{}\mathpunct{},+\infty\mathclose{)}$ and such that
@@
$\frac{e^{-X_{j}\left(b,u,q\right)\cosh t}\left(1+X_{j}\left(b,u,q\right)\cosh t+e^{-X_{j}\left(b,u,q\right)\cosh t}\right)}{\left(1+e^{-X_{j}\left(b,u,q\right)\cosh t}\right)^{2}}\\ \leq e^{-X_{j}\left(b,u,q\right)\cosh t}\left(1+X_{j}\left(b,u,q\right)\cosh t\right)\\ \leq e^{-X_{j}\left(b,u,q\right)\cosh t}+e^{-\frac{1}{2}X_{j}\left(b,u,q\right)\cosh t}\\ \leq 2e^{-\frac{1}{2}X_{j}\left(b,u,q\right)\cosh t}\\ \leq 2e^{-\frac{1}{2}X_{j}\left(b,u,q\right)\left(1+\frac{t^{2}}{2}\right)},$
@@
$\frac{2}{\beta}\int_{0}^{1}u\left(1-u\right)\int_{0}^{\beta}e^{-\frac{1}{2}X_{j}\left(b,u,q\right)}\int_{0}^{\infty}e^{-\frac{t^{2}}{4}X_{j}\left(b,u,q\right)}\mathop{}\mathopen{}\mathrm{d}t\mathop{}\mathopen{}\mathrm{d}b\mathop{}\mathopen{}\mathrm{d}u\\ =\frac{\sqrt{\pi}}{3\sqrt{m_{j}}\beta}\int_{0}^{\beta}\frac{1}{\sqrt{X_{j}\left(b,u,q\right)}}e^{-\frac{1}{2}X_{j}\left(b,u,q\right)}\mathop{}\mathopen{}\mathrm{d}b\\ \leq\frac{2\sqrt{\pi}}{3\sqrt{m_{j}}}\frac{1}{\sqrt{\beta}}.$
In particular, the previous calculations show that the function $M^{T}$ is well defined and bounded for any fixed $\beta\in\mathopen{(}0\mathclose{}\mathpunct{},+\infty\mathclose{)}$. Now, in order to conclude the proof, let us recall that
@@
and $G\left(q,\beta\right)\geq 0$ by [34, Lemma 11]. Indeed,